# From Light-Cone to Supersonic Propagation of Correlations by Competing Short- and Long-Range Couplings

### v1

@@ D Truncated Wigner approximation
$+i\Omega(\hat{a}+\hat{a}^{\dagger})(-1)^{j}\hat{b}_{j}$ (D.1)
$\frac{d\hat{a}}{dt}=$ $-\left(i\delta+\frac{\Gamma}{2}\right)\hat{a}+\sqrt{\Gamma}\hat{a}_{\mathrm{in}}(t)$ (D.2)
@@
$\frac{d\beta_{j}}{dt}=$ $i\frac{J}{\hbar}\sum_{l\in n(j)}\beta_{l}-i\frac{U}{\hbar}\left(|\beta_{j}|^{2}-\frac{1}{2}\right)\beta_{j}$ (D.3)
@@
$\frac{d\alpha}{dt}=$ $-\left(i\delta+\frac{\Gamma}{2}\right)\alpha+\sqrt{\Gamma}\mathcal{F}(t)$ (D.4)
@@
with $\mathcal{F}(t)=[\mathcal{F}_{r}(t)+i\mathcal{F}_{i}(t)]/2$, $\langle\mathcal{F}_{r}\rangle=0=\langle\mathcal{F}_{i}\rangle$, $\langle\mathcal{F}_{r}(t^{\prime})\mathcal{F}_{i}(t)\rangle=0$, and $\langle\mathcal{F}_{i}(t^{\prime})\mathcal{F}_{i}(t)\rangle=\delta(t-t^{\prime})$, $\langle\mathcal{F}_{r}(t^{\prime})\mathcal{F}_{r}(t)\rangle=\delta(t-t^{\prime})$. Equations (D.3) and (D.4) are the dynamical equations we simulate in our truncated Wigner approach with a given initial condition.
@@ 1 Initial conditions
To model the initial condition of the atomic Fock state and the empty cavity states used in the simulations of the main text, we initialize the cavity in the vacuum state and sample $\alpha^{(r)}(t=0)$ and $\alpha^{(i)}(t=0)$ with $\alpha(t=0)=\alpha^{(r)}(0)+i\alpha^{(i)}(0)$ as Gaussian random variables with $\langle\alpha^{(r)}(0)\rangle=0=\langle\alpha^{(i)}(0)\rangle$, $\langle\alpha^{(r)}\alpha^{(i)}\rangle=0$, and $\langle\alpha^{(r)}(0)\alpha^{(r)}(0)\rangle=1/4=\langle\alpha^{(i)}(0)\alpha^{(i)}(0)\rangle$. For the atomic state we assume that each mode indexed by $j$ is in a Fock state with occupation $n_{j}$. For this we employ a sampling method which was previously derived in Refs. Gardiner et al. (2002); Olsen et al. (2004); Nagao et al. (2019). We sample $\beta_{j}(0)=[p_{j}+q_{j}\eta_{j}]e^{i\varphi_{j}}$ where $\eta_{j}$ is a Gaussian random variable with mean equal zero and with variance $\langle\eta_{j}^{2}\rangle=1$ and $\varphi_{j}$ is a uniform random variable in the interval $[0,2\pi)$. The values of $p_{j}$ and $q_{j}$ are determined by the occupation $n_{j}$ with
$p_{j}=$ $\frac{\sqrt{2n_{j}+1+2\sqrt{n_{j}^{2}+n_{j}}}}{2},$ (D.5)
$q_{j}=$ $\frac{1}{4p_{j}}.$ (D.6)
@@
$\langle\beta_{j}^{*}(0)\beta_{j}(0)\rangle=$ $n_{j}+\frac{1}{2}$ (D.7)
@@
$\langle\beta_{j}^{*}(0)\beta_{j}^{*}(0)\beta_{j}(0)\beta_{j}(0)\rangle=$ $\left(n_{j}+\frac{1}{2}\right)^{2}+\frac{1}{4}.$ (D.8)
@@ 2 Technical details on the simulations
For the simulation of the stochastic differential equation we have used a Runge-Kutta method of 4th order with an integration time step $dt=0.5\times 10^{-2}/\Gamma$. All results shown in the main text are obtained from $M=10^{6}$ noise initializations. For each of the initialization that we index in the following by $m=1,...,M$ we obtain fluctuating densities
$n^{(m)}_{j}(t)=|\beta^{(m)}_{j}(t)|^{2}-\frac{1}{2}.$ (D.9)
@@
$\mathcal{C}_{nn}(d,t)=$ $\frac{1}{M}\sum_{m=1}^{M}n^{(m)}_{j_{0}}(t)\frac{n^{(m)}_{j_{0}+d}(t)+n^{(m)}_{j_{0}-d}(t)}{2}$ (D.10)
$-\frac{1}{M^{2}}\sum_{m,m^{\prime}=1}^{M}n^{(m^{\prime})}_{j_{0}}(t)\frac{n^{(m)}_{j_{0}+d}(t)+n^{(m)}_{j_{0}-d}(t)}{2}$ (D.11)
where $j_{0}=(L+1)/2$ is the center and the division by $2$ comes from the fact that there are two sites with the same distance in 1D. The result of this calculation is shown in Fig. 3(c) in the main text.
@@
$\mathcal{C}_{nn}(d,t)=$ $\frac{1}{M}\sum_{m=1}^{M}n^{(m)}_{\mathbf{j}_{0}}(t)\sum_{{\bf j}}\frac{n^{(m)}_{{\bf j}}(t)}{\mathcal{N}_{d}}\delta_{d({\bf j}),d}$ (D.12)

### v2

@@ E Truncated Wigner approximation (TWA)
$+i\Omega(\hat{a}+\hat{a}^{\dagger})(-1)^{j}\hat{b}_{j}$ (E.1)
$\frac{d\hat{a}}{dt}=$ $-\left(i\delta+\frac{\Gamma}{2}\right)\hat{a}+\sqrt{\Gamma}\hat{a}_{\mathrm{in}}(t)$ (E.2)
@@
$\frac{d\beta_{j}}{dt}=$ $i\frac{J}{\hbar}\sum_{l\in n(j)}\beta_{l}-i\frac{U}{\hbar}\left(|\beta_{j}|^{2}-\frac{1}{2}\right)\beta_{j}$ (E.3)
@@
$\frac{d\alpha}{dt}=$ $-\left(i\delta+\frac{\Gamma}{2}\right)\alpha+\sqrt{\Gamma}\mathcal{F}(t)$ (E.4)
@@
with $\mathcal{F}(t)=[\mathcal{F}_{r}(t)+i\mathcal{F}_{i}(t)]/2$, $\langle\mathcal{F}_{r}\rangle=0=\langle\mathcal{F}_{i}\rangle$, $\langle\mathcal{F}_{r}(t^{\prime})\mathcal{F}_{i}(t)\rangle=0$, and $\langle\mathcal{F}_{i}(t^{\prime})\mathcal{F}_{i}(t)\rangle=\delta(t-t^{\prime})$, $\langle\mathcal{F}_{r}(t^{\prime})\mathcal{F}_{r}(t)\rangle=\delta(t-t^{\prime})$. Equations (E.3) and (E.4) are the dynamical equations we simulate in our truncated Wigner approach with a given initial condition.
@@ 1 Initial conditions
To model the initial condition of the atomic Fock state and the empty cavity states used in the simulations of the main text, we initialize the cavity in the vacuum state and sample $\alpha^{(r)}(t=0)$ and $\alpha^{(i)}(t=0)$ with $\alpha(t=0)=\alpha^{(r)}(0)+i\alpha^{(i)}(0)$ as Gaussian random variables with $\langle\alpha^{(r)}(0)\rangle=0=\langle\alpha^{(i)}(0)\rangle$, $\langle\alpha^{(r)}\alpha^{(i)}\rangle=0$, and $\langle\alpha^{(r)}(0)\alpha^{(r)}(0)\rangle=1/4=\langle\alpha^{(i)}(0)\alpha^{(i)}(0)\rangle$. For the atomic state we assume that each mode indexed by $j$ is in a Fock state with occupation $n_{j}$. For this we employ a sampling method which was previously derived in Refs. Gardiner et al. (2002); Olsen et al. (2004); Nagao et al. (2019). We sample $\beta_{j}(0)=[p_{j}+q_{j}\eta_{j}]e^{i\varphi_{j}}$ where $\eta_{j}$ is a Gaussian random variable with mean equal zero and with variance $\langle\eta_{j}^{2}\rangle=1$ and $\varphi_{j}$ is a uniform random variable in the interval $0,2\pi)$. The values of $p_{j}$ and $q_{j}$ are determined by the occupation $n_{j}$ with
$p_{j}=$ $\frac{\sqrt{2n_{j}+1+2\sqrt{n_{j}^{2}+n_{j}}}}{2},$ (E.5)
$q_{j}=$ $\frac{1}{4p_{j}}.$ (E.6)
@@
$\langle\beta_{j}^{*}(0)\beta_{j}(0)\rangle=$ $n_{j}+\frac{1}{2}$ (E.7)
@@
$\langle\beta_{j}^{*}(0)\beta_{j}^{*}(0)\beta_{j}(0)\beta_{j}(0)\rangle=$ $\left(n_{j}+\frac{1}{2}\right)^{2}+\frac{1}{4}.$ (E.8)
@@ 2 Technical details on the simulations
For the simulation of the stochastic differential equation we have used a Runge-Kutta method of 4th order with an integration time step $dt=0.5\times 10^{-2}/\Gamma$. All results shown in the main text are obtained from $M=10^{6}$ noise initializations. For each of the initialization that we index in the following by m=1,,Mm=1,...,M we obtain fluctuating densities
$n^{(m)}_{j}(t)=|\beta^{(m)}_{j}(t)|^{2}-\frac{1}{2}.$ (E.9)
@@
$\mathcal{C}_{nn}(d,t)=$ $\frac{1}{M}\sum_{m=1}^{M}n^{(m)}_{j_{0}}(t)\frac{n^{(m)}_{j_{0}+d}(t)+n^{(m)}_{j_{0}-d}(t)}{2}$ (E.10)
$-\frac{1}{M^{2}}\sum_{m,m^{\prime}=1}^{M}n^{(m^{\prime})}_{j_{0}}(t)\frac{n^{(m)}_{j_{0}+d}(t)+n^{(m)}_{j_{0}-d}(t)}{2}$ (E.11)
where $j_{0}=(L+1)/2$ is the center and the division by 22 comes from the fact that there are two sites with the same distance in 1D. The result of this calculation is shown in Fig. 3(c) in the main text.
@@
$\mathcal{C}_{nn}(d,t)=$ $\frac{1}{M}\sum_{m=1}^{M}n^{(m)}_{\mathbf{j}_{0}}(t)\sum_{{\bf j}}\frac{n^{(m)}_{{\bf j}}(t)}{\mathcal{N}_{d}}\delta_{d({\bf j}),d}$ (E.12)
@@
F Local dissipation
Figure F1: Correlations $\mathcal{C}_{nn}(d,t)$ calculated from TWA simulations for (a) vanishing local dephasing $\gamma=0$ and (b) $\hbar\gamma/J=1$ in 1D for a lattice with $L=21$ sites. We simulated Eq. (E.3) added Eq. (F.3) and Eq. (E.4) with $U/J=0.1$, $\hbar\delta/J=2$, $\hbar\Gamma/J=1$, $\sqrt{N}\hbar\Omega/J=4$. The atoms are initialized with alternating densities of $n=10$ and $n=11$ bosons corresponding to total atom number $N=220$ and the cavity is initialized in the vacuum state.
In this section we show briefly that supersonic spreading can also be observed in presence of local atomic dissipation. To study the effect of local dissipation we use the TWA simulation introduced in the previous section.
Local dissipation is added to our previous model in form of local dephasing
$\mathcal{L}_{j}\hat{\rho}$= $-\frac{\gamma}{2}(\hat{n}_{j}^{2}\hat{\rho}+\hat{\rho}\hat{n}_{j}^{2}-2\hat{n}_{j}\hat{\rho}\hat{n}_{j})$ (F.1)
with rate $\gamma$ and such that the master equation governing the dynamics of the density operator $\hat{\rho}$ reads
[…]
$+$ $\sum_{j}\mathcal{L}_{j}\hat{\rho}$. (F.2)
Employing symmetric ordering we can derive the dynamics for the complex-valued classical fields $\beta_{j}=\beta_{j}^{(r)}+i\beta_{j}^{(i)}$. The local dephasing term results in a term which needs to be added to Eq. (E.3) which is
$\left(\frac{d\beta_{j}}{dt}\right)_{\mathrm{local}}=-\frac{\gamma}{2}\beta_{j}-i\sqrt{\gamma}\beta_{j}W_{j}.$ (F.3)
Here we introduced local noise terms $W_{j}$ that obey $\langle W_{j}(t)\rangle=0$ and $\langle W_{j}(t)W_{i}(t^{\prime})\rangle=\delta_{ij}\delta(t-t^{\prime}).$
We simulate the corresponding coupled stochastic differential equation of the complex-valued fields $\alpha$ and $\beta_{j}$ for the same parameters as in Fig. 3(c) in the main text. In Fig. F1(a) and (b) we show the results of the simulations for vanishing $\gamma=0$ and for non-vanishing $\hbar\gamma/J=1$, respectively. We find slight modifications in the magnitude of the correlations. In particular, the value of $\mathcal{C}_{nn}$ seems to be slightly smaller for $\hbar\gamma/J=1$ than it is for $\gamma=0$. In general, however, the shape of the correlation spreading is very similar for both cases. We highlight that we observe the light-cone like spreading for short times and the simultaneous supersonic spreading afterwards. This finding confirms that the features that are discussed in the paper are robust against (weak) local dephasing terms.